A liter of cream contains 9.5 percent butterfat. How much skim milk containing 5.5 percent butterfat must be added to the mixture to obtain one liter of mixture with 6.5 percent butterfat?
Answer by ankor@dixie-net.com(22740)

A liter of cream contains 9.5 percent butterfat. How much skim milk containing 5.5 percent butterfat must be added to the mixture to obtain one liter of mixture with 6.5 percent butterfat?
:
Let x = amt (in liters)of skim milk to be added
:
Write it as a per-cent butter fat equation, you don't have to convert the % to decimals here, unless you just want to.
:
5.5x + 9.5(1) = 6.5(x+1)
5.5x + 9.5 = 6.5x + 6.5
5.5x - 6.5x = 6.5 - 9.5
-1x = -3
x = + 3 liters of skim milk
:
Check it using decimals:
.055(3) + .095(1) = .065(3+1)
.165 + .095 = .260